A stock solution of 1000 ppm (w/V) Cu(II) was prepared from Cu(NO3)2.  

1. Calculate the mass of the solid needed to prepare 1.00 L.    

           

2.    Determine the volumes needed to prepare 50.0 mL calibration standards with concentrations

           of 5.00 ppm; 10.0 ppm and 25.0 ppm, respectively.     

       

1
Expert's answer
2021-01-18T01:31:29-0500

1. 1000 ppm means that there is 1g of Cu(II) in 1000 ml.

w(Cu)=Ar(Cu)/Mr(Cu(NO3)2)

Mr(Cu(NO3)2)=64+2×14+6×16=188(g/mol)

w(Cu)=64/188=0.34

m(Cu(NO3)2)=m(Cu)/0.34= 1/0.34=2.94(g)


2. C1V1=C2V2

V1=C2V2/C1

1) V1=5×50/1000=0.25(ml)

2)V1=10×50/1000=0.5(ml)

3)V1=25×50/1000=1.25(ml)